A heat engine operates between a high temperature of about 600?C and a low temperature of about 300?C. What is the maximum theoretical efficiency for this engine?

A heat engine operates between a high temperature of about 600 and a low temperature of about 300. What is the maximum theoretical efficiency for this engine?

A) =100%.
B) ?66%.
C) ?50%.
D) ?34%.
E) Cannot be determined from the given information.
Answer #1

efficiency = 1 - (573 / 873) = 34%

D) 34%
Answer #2

ans is D

efficiency = 1 - (TH/TC)

34